# Airborne Sense and Detect of Drones using Deep Learning and LiDAR Point Clouds

### v2

@@ Abstract
The safe operation of drone swarms beyond visual line of sight requires multiple safeguards to mitigate the risk of collision between drones flying in close-proximity scenarios. Cooperative navigation and flight coordination strategies that rely on pre-planned trajectories, constant network connectivity and reliable Global Navigation Satellite System (GNSS) positioning are brittle to failure. Drone embedded sense and detect offers a comprehensive mode of separation between drones for deconfliction and collision avoidance. This paper presents the first airborne LiDAR based solution for drone-swarm detection and localization using 3D deep learning model. It adapts an existing deep learning neural network to the air-to-air drone scenario by expanding the scan space vertically. A new sparse convolution is proposed and applied to accelerate the backbone layer, which is the most time-consuming part of the neural network. To collect training data of safety critical, close-proximity multi-drone operations, a scenario Digital Twin is used to augment real datasets with high fidelity synthetic data. The trained model achieves over 80% recall and 96% precision when tested on real-world datasets. By incorporating a tracking-by-detection algorithm the system can reliably monitor the separation distance of multiple drones in challenging environments.
@@ 1 Introduction
Sense and detect, which involve identifying and responding to potential conflicts, are crucial for estimating risks in dynamic drone swarm environments [1]. As drone use cases expand, it becomes essential to validate technologies that can take on the safety-critical role of onboard pilots by monitoring surrounding airspace for hazards, including air and ground collision risks. Current drone swarm management systems rely primarily on trajectory-based flight planning, where predefined flight paths guide drones to avoid conflicts. Simultaneous Localization and Mapping (SLAM), which enables drones to build a map of their environment and track their position within it, along with anti-spoofing techniques that protect against falsified GPS signals, have improved in-flight positional estimation and broadcast. However, tactical deconfliction—the process of managing and resolving potential flight path conflicts—still depends on centralized decision-making and continuous connectivity to a command and control (C2) center. This centralized system is vulnerable to network failures, and it relies on cooperative behavior from all airborne vehicles. Faulty or uncooperative traffic can enter the flight space undetected, posing risks. To safeguard users and mitigate air-to-air collision risks, onboard drone-embedded sense and detect capabilities, where drones autonomously perceive and avoid conflicts, are necessary.
This paper addresses the challenge of monitoring a drone swarm while they perform a collaborative mission, such as a maintenance-inspection survey. The integrated system can effectively monitor the separation distance between multiple drones in challenging environments and alert for potential air-air collisions between drones. The paper describes the first drone-embedded system that uses point-based artificial intelligence for detection and localization of other drones. We adapt an existing detection model, originally designed for ground applications such as autonomous vehicles, to an air-to-air drone scenario by expanding the scan space vertically. To reduce the latency of the detection model, we introduce a new sparse convolution algorithm that leverages the sparsity of LiDAR data. This algorithm is applied to accelerate the backbone layer of the model, which is the most time-consuming component of the detection process.
Currently, drones present challenges in the AI domain, characterized by a lack of high-quality data for training and validating hazard perception algorithms, coupled with an absence of data sharing mechanisms that can handle sensitive and privacy constrained hazard observation data. We develope a blended reality data augmentation strategy that outperforms conventional data augmentation strategies by adapting a scenario Digital Twin of the operating scenario and equipping the virtual environment with high fidelity drone and LiDAR simulators. Our results demonstrate that detection models trained on high-fidelity synthetic augmentation datasets outperform those trained on conventional augmentation datasets.
Our contributions are:
@@
    We propose the first airborne LiDAR based solution for drone-swarm detection and localization using 3D deep learning, which considers critical factors of real-time sense and detect of drones for safety scenarios. We adapt an existing detection model for an air-to-air drone scenario by expanding the scan space vertically.
@@
    We propose a novel sparse convolution algorithm to accelerate the backbone layer of the detection model, achieving a speedup factor of 2.3.
@@
    To further improve detection accuracy, we employ a synthetic data augmentation strategy for training data generation. The model trained on this dataset achieves the highest performance with 80% recall and a 96% precision. The model, trained on real-world data, achieves a recall of 62% and a precision of 82.6%.
@@ 2 Related Work
In the context of drone conflict avoidance, sense and detect can be defined as follows: Sense refers to the methods employed by a drone to actively gather information about its surrounding environment. These methods typically involve onboard sensor technologies, such as cameras, LiDAR, and other advanced sensing systems. Detect involves analyzing the data collected by the drone’s sensors. The analysis aims to identify the presence of threats and hazards within the drone’s operational space and assess the potential and severity of these threats to the drone’s safe operation.
@@ 2.1 Detection and Tracking of Drones
2.1.1 Data collection using various sensors
@@
An exploration [2] is conducted to examine the distinctive marks of small Unmanned Aerial Vehicles (UAVs) utilizing a variety of sensor technologies. These technologies include acoustical antennas, passive and active optical imaging tools, as well as compact FMCW RADAR systems. As described in [3], acoustic-based detection of UAVs utilizing various machine learning algorithms offers a promising approach for UAV identification. However, this method inherently faces limitations in precise localization in 3D space which is critical requirement in our use case. While [4] presents a method for single-drone detection using passive radio frequency (RF) signals, this technique is less effective for drone swarms. The presence of multiple drones operating in the same area (as in our use case) elevates the overall background RF noise level, hindering the ability of RF-based detection systems to isolate and identify the weaker signals from individual drones within the swarm.
@@
Radar technology [5][6][7] presents a valuable approach for drone detection, though its effectiveness is hampered by the low radar cross-section (RCS) of drones. This limitation can lead to missed detections, particularly for smaller drones, significantly restricting its use in safety-critical applications such as our use case involving drone swarm monitoring. Several camera-based detect and track of small UAVs are presented in [8][9][10][11]. Camera-based UAV detection offers an efficient solution for verification and classification of drones, however, their performance is significantly impacted by environmental factors like lighting conditions and weather. Additionally, cameras inherently provide only 2D information, lacking the crucial depth data required for precise 3D localization, which is essential for our needs.
@@
A complete chain of detection, tracking, and classification of small flying objects in real-time is presented in [12] . This is achieved through the use of a mobile multi-sensor platform equipped with two 360 LiDAR scanners and pan-and-tilt cameras in the visible and thermal infrared (IR) spectrum. The approach involves the initiall detect and track of flying objects in 3D LiDAR data. After localizing, the cameras are automatically directed to the object’s position, and each sensor records a 2D image. A convolutional neural network (CNN) is then used to identify the region of interest (ROI) and classify the object into one of eight different types of UAVs and birds. Clearly this two-stage detection and classification approach will introduce higher latency. Alongside latency issues, deploying this system on a drone poses further challenges due to the weight and power limitations inherent in drone platforms.
@@
As UAVs have gained popularity, the attention towards cooperative UAV detection and tracking has intensified. To improve cooperative UAV navigation in GNSS-challenging conditions, the approach presented in [13] combines camera based measurements with navigation data from other aircraft flying under better GNSS coverage. Clearly, this approach is not independent of external aiding systems. External aiding systems, such as GNSS, can introduce multipath errors caused by signals reflecting off buildings, trees, the ground, and other surfaces.
@@
This paper presents the first airborne LiDAR based solution for drone detection using point-cloud based 3D deep learning network. This approach represents an advancement in drone detection technology, as it satisfies different requirements for real-time sensing and detection of drones in safety critical scenarios.
@@ 2.1.2 Deep learning on points clouds
Recently deep learning on point clouds has become a prevalent AI technique in autonomous driving. Numerous methods have been proposed to address different problems in this area. A comprehensive review of deep learning approaches on point clouds is presented in [14]. As a pioneering work on point clouds classification and segmentation, PointNet [15] and its successor PointNet++ [16] has inspired many others works in this area.
@@
VoxelNet [17] organizes point clouds into evenly spaced 3D voxels and applies a PointNet to encode point-wise features within each voxel. A 3D convolutional middle layer and a region proposal network are then applied to generate 3D object proposals in the form of 3D bounding boxes. The computational bottleneck resides in the 3D convolution layer. SECOND [18] improved VoxelNet by applying sparse convolution and achieved an inference speed of 30.4 FPS [19] on a NVIDIA RTX 2080 Ti GPU, but the 3D convolutions remain a bottleneck. PointPillars [20] enables end-to-end learning with only 2D convolutional layers. It divides the x-y coordinate plane into evenly spaced grid cells and organizes point clouds into the corresponding cell to create ”pillars”. A simplified PointNet is applied the points within pillars. The backbone network consisting of multiple 2D convolutions is then applied to pseudo-image of features generated from the processed pillars. A detection head [21] is then used to generate 3D bounding box for objects in point cloud data. The inference of PointPillars can run at a speed of 42.4 FPS. Therefore it is highly suitable for applications requiring real time processing.
@@
PV-RCNN [22] extends SECOND by adding a keypoints branch to preserve 3D structural information. Its backbone network takes the raw point cloud data as input to extract point-wise features. PV-RCNN can achieves a higher accuracy but its inference speed is only 8.9 FPS, which falls below real-time requirements. Voxel-RCNN [19] further improves accuracy by leveraging a 3D network for feature extraction. It then applies a 2D backbone network and Region Proposal Network (RPN) to BEV representation of the 3D feature volumes. Its inference achieves the speed of 25.2 FPS.
@@
While there are various 3D object detection networks that can achieve higher detection performance, we selected PointPillars for our work due to its real-time inference capability; it consistently achieves the highest FPS compared to other detection models [19]. This is a critical consideration for our use-case application, which is safety orientated and requires the ability to process incoming point clouds in real time. While higher detection performance is desirable, it often comes at the cost of increased computation time and resource requirements, which can limit the feasibility of deploying the network in real-world application scenarios. We chose PointPillars as a balance between detection performance and real-time inference capability to meet the requirements of our drone detection applications.
@@
Instead of purely using the original PointPillars network, we modify it for an air-to-air drone scenario by extending the scan space vertically. To reduce inference latency, we propose a new sparse convolution algorithm to accelerate the backbone layer, which is the most time-consuming part of PointPillars [23].
@@ 2.2 Latency Reduction
While increasing the number of point clouds can improve detection accuracy, minimizing latency is crucial for ensuring real-time hazard response. The inherent sparsity of LiDAR data can be utilized to reduce latency. The study in [24] describes the spatial sparsity of feature maps and how sparsity applies to improve performance of different applications. A sparse convolutional neural network that utilizes sparse matrix multiplication is presented in [25]. An efficient sparse convolutional operation called submanifold sparse convolution is introduced in [26], in which the output will be computed if and only if the corresponding site in the input is active.
Matrix multiplication is an efficient method for implementing spatial convolution on GPU, as demonstrated by [27]. This requires transforming the input image into matrices that are suitable for fast multiplication. This transformation can be achieved by lowering the input data with duplication. The GPU implementation of the proposed method in this paper is also based on fast GPU matirix multiplication, but without data lowering.
A sparse convolution with rule book is presented in [18], where the rule book is used to map between input and output. The work presented in [28] generalizes convolution, including sparse convolution, to handle any discrete convolutions in high-dimensional spaces. A highly efficient Sparse Kernel Generator is developed in [29], which produces high performance sparse point cloud convolution kernels at a fraction of the engineering cost of the current state-of-the-art.
Most current methods implement convolution as a gathering operation, where neighboring pixels are multiplied by corresponding filter elements, and the results are summed to update the center pixel. Sparse convolution implementations typically follow this approach as well. To accelerate GPU performance, a hash table or rule book is often used to map positions between the input and output. In this paper, we propose an alternative approach to implementing sparse convolution using a scattering operation. This method eliminates the need for a hash table or rule book, improving efficiency while simplifying the implementation.
@@ 2.3 Training Data Generation
Within the context of autonomous driving, there are different dataset such as KITTI [30], Waymo Open [31] and nuScenes [32] and others which are available online, and include LiDAR as part of the dataset. These datasets contain a collection of diverse urban scenes and dense position and class annotations. These datasets provide us with valuable resources for validating autonomous driving algorithms, enabling the development of more accurate and robust perception systems.
@@
LiDAR is one of the most important sensors for autonomous driving and many works have focused on the simulation of LiDAR. In [38][39], a LiDAR simulator which can augment real point clouds with synthetic objects is implemented. The real point cloud (usually a background) is generated from real LiDAR scanners mounted on top of a vehicle sweeping real world surroundings. The point cloud of a synthetic object is obtained by scanning handcraft CAD models with virtual LiDAR. These works demonstrate the importance of position and orientation of scanning objects for detection and segmentation. A LiDAR simulation approach described in [40] leverages real-world data to enhance the realism of the simulated scenes. This approach results in more diversity and realism in the simulated scenes. It also describes a learning system that models the residual physics that are not captured by graphics rendering, which further boosts the realism of the simulation. For a drone usage scenario, the study in [41] employs LiDAR simulation to evaluate the performance of detect-and-avoid systems for unmanned aerial systems (UAS).
In this work, we enrich the AI training dataset by using a Digital Twin to accurately simulate a LiDAR sensor and augment the simulated point clouds within a real-world background. Unlike conventional augmentation techniques, our approach reflects the actual position and orientation of the scanned object.

### v3

# Real-Time Sense and Detect of Drones Using Deep Learning and Airborne LiDAR

@@ Abstract
The safe operation of drone swarms beyond visual line of sight requires multiple safeguards to mitigate the risk of collision between drones flying in close-proximity scenarios. Cooperative navigation and flight coordination strategies that rely on pre-planned trajectories, constant network connectivity and reliable Global Navigation Satellite System (GNSS) positioning are brittle to failure. Drone embedded sense and detect offers a comprehensive mode of separation between drones for deconfliction and collision avoidance. This paper presents the first airborne LiDAR-based solution for drone-swarm detection and localization using a 3D deep learning model. It adapts an existing deep learning neural network to the air-to-air drone scenario by expanding the scan space vertically. A new sparse convolution is proposed and applied to accelerate the backbone layer, which is the most time-consuming part of the neural network. To collect training data of safety critical, close-proximity multi-drone operations, a scenario Digital Twin is used to augment real datasets with high-fidelity synthetic data. The trained model achieves over 80% recall and 96% precision when tested on real-world datasets.
@@ 1 Introduction
Sense and detect algorithms which involve identifying potential conflicts, are crucial for assessing risks in dynamic drone swarm environments [1]. In the context of drone conflict avoidance, sense and detect can be defined as follows: Sense refers to the methods employed by a drone to actively gather information about its surrounding environment. These methods typically involve onboard sensor technologies, such as cameras, LiDAR, and other advanced sensing systems. Detect involves analyzing the data collected by the drone’s sensors. The analysis aims to identify the presence of threats and hazards within the drone’s operational space and assess the potential and severity of these threats to the drone’s safe operation. Together, these functions allow the drone to recognize obstacles and assess risks in real time, forming the foundation for effective avoidance strategies.
@@
As drone use cases expand, it becomes essential to validate technologies that can take on the safety-critical role of onboard pilots by monitoring surrounding airspace for hazards, including air and ground collision risks. Current drone swarm management systems rely primarily on trajectory-based flight planning, where predefined flight paths guide drones to avoid conflicts. Simultaneous Localization and Mapping (SLAM), which enables drones to build a map of their environment and track their position within it, along with anti-spoofing techniques that protect against falsified GPS signals, have improved in-flight positional estimation and robustness. However, tactical deconfliction—the process of managing and resolving potential flight path conflicts—still depends on centralized decision-making and continuous connectivity to a command and control (C2) center. This centralized system is vulnerable to network failures, and it relies on cooperative behavior from all drones. To safeguard users and mitigate air-to-air collision risks, onboard drone-embedded sense and detect capabilities, where drones autonomously perceive and avoid conflicts, are necessary.
While ensuring safe separation in drone swarms is critical for collision avoidance, these capabilities also enable novel applications beyond airspace management. One such application is the inspection of critical infrastructure, where autonomous drone swarms can significantly enhance efficiency and coverage [2][3][4]. Multiple civil engineering reports state that 1 in 10 bridges are at risk of failure [5][6], underscoring the need for advanced inspection solutions. Our research demonstrates that drone swarms can enable continuous large-scale infrastructure inspection and monitoring in complex transport system environments, reducing operational downtime and enhancing coverage efficiency. In our drone swarm framework, a monitor drone utilizes LiDAR-based 3D object detection to track and coordinate the swarm, mitigating collision risks caused by GNSS drift [7] and localization errors. The approach improves the safety, reliability, and efficiency of inspections. This paper focuses on the 3D point cloud detection model. Our contributions are:
@@
    We present a real-time airborne LiDAR-based solution for drone-swarm detection and localization using 3D deep learning, designed to operate in complex transportation infrastructure environments. The model introduces altitude-stratified anchor boxes to improve detection robustness for airborne objects at varying flight levels, enabling effective sense-and-detect capabilities in high-risk swarm scenarios.
@@
    A novel sparse convolution algorithm is developed to achieve real-time inference on airborne embedded platforms, leveraging a scattering-based implementation that eliminates the need for hash tables or rule books. This optimization significantly reduces inference latency, enabling deployment on edge hardware while maintaining detection accuracy. The algorithm accelerates the detection model, achieving a 2.3× speedup.
@@
    A scenario-driven Digital Twin is used to generate high-fidelity synthetic training data, simulating LiDAR scans of drones in diverse safety-critical scenarios. This blended reality augmentation approach enhances model performance, achieving 80% recall and 96% precision, compared to 62% recall and 82.6% precision when trained solely on real-world data.
@@ 2 Related Work
In drone conflict avoidance, a drone gathers environmental data using onboard sensors such as cameras and LiDAR, then analyzes this data to identify potential threats and assess their severity to ensure safe operation.
@@ 2.1 Detection and Tracking of Drones
2.1.1 Multisensor Approaches for Drone Detection
@@
A survey [8] is conducted to examine the distinctive marks of small drones utilizing a variety of sensor technologies. These technologies include acoustical antennas, passive and active optical imaging tools, as well as compact FMCW RADAR systems. As described in [9], acoustic-based detection of drones utilizing various machine learning algorithms offers a promising approach for drone identification. However, this method inherently faces limitations in precise localization in 3D space. While [10] presents a method for single-drone detection using passive radio frequency (RF) signals, this technique is less effective for drone swarms. The presence of multiple drones operating in the same area elevates the overall background RF noise level, hindering the ability of RF-based detection systems to isolate and identify the weaker signals from individual drones within the swarm.
@@
RADAR technology [11][12][13] presents a valuable approach for drone detection, though its effectiveness is hampered by the low radar cross-section (RCS) of drones. This limitation can lead to missed detections, particularly for smaller drones, significantly restricting its use in safety-critical applications of drone swarm monitoring. Several camera-based detect and track of small drones are presented in [14][15][16][17]. Camera-based drone detection offers an efficient solution for detection and classification of drones, however, their performance is significantly impacted by environmental factors like lighting conditions and weather. Additionally, cameras inherently provide only 2D information, lacking the crucial depth data required for precise 3D localization. While stereo vision [18] can offer depth estimates, accurate depth for larger distances requires prohibitively long stereo baselines, which can be challenging for our needs.
@@
A complete chain of detection, tracking, and classification of small flying objects in real-time is presented in [19]. This is achieved through the use of a mobile multi-sensor platform equipped with two 360 LiDAR scanners and pan-and-tilt cameras in the visible and thermal infrared (IR) spectrum. The approach involves the initial detect and track of flying objects in 3D LiDAR data. After localizing, the cameras are automatically directed to the object’s position, and each sensor records a 2D image. A convolutional neural network (CNN) is then used to identify the region of interest (ROI) and classify the object into one of eight different types of drones and birds. Clearly this two-stage detection and classification approach will introduce higher latency. Alongside latency issues, deploying this system on a drone poses further challenges due to the weight and power limitations inherent in drone platforms.
@@
As drones have gained popularity, the attention towards cooperative drone detection and tracking has intensified. To improve cooperative drone navigation in challenging GNSS conditions, the approach presented in [20] combines camera based measurements with navigation data from other aircraft flying under better GNSS coverage. Clearly, this approach is not independent of external positioning systems such as GNSS which can introduce multipath errors caused by signals reflecting off buildings, trees, the ground, and other surfaces.
@@
This paper presents an airborne LiDAR based solution for drone detection using a point cloud-based 3D deep learning network. This approach represents an advancement in drone detection technology, as it satisfies different requirements for real-time sense and avoid of drones in safety critical scenarios. These requirements include low latency, high detection performance, 3D localization capability, independence from external positioning systems, and the ability to operate within the strict power and weight limitations of drones.
@@ 2.1.2 Deep learning on points clouds
Recently deep learning on point clouds has become a prevalent AI technique in autonomous driving. Numerous methods have been proposed to address different problems in this area. A comprehensive review of deep learning approaches on point clouds is presented in [21]. As a pioneering work on point clouds classification and segmentation, PointNet [22] and its successor PointNet++ [23] have inspired many other works in this area.
@@
VoxelNet [24] organizes point clouds into evenly spaced 3D voxels and applies a PointNet to encode point-wise features within each voxel. A 3D convolutional middle layer and a region proposal network are then applied to generate 3D object proposals in the form of 3D bounding boxes. The computational bottleneck resides in the 3D convolution layer. SECOND [25] improved VoxelNet by applying sparse convolution and achieved an inference speed of 30.4 FPS [26] on a NVIDIA RTX 2080 Ti GPU, but the 3D convolutions remain a bottleneck. PointPillars [27] enables end-to-end learning with only 2D convolutional layers. It divides the x-y coordinate plane into evenly spaced grid cells and organizes point clouds into the corresponding cell to create ”pillars”. A simplified PointNet is applied to the points within pillars. A backbone network consisting of multiple 2D convolutions is then applied to pseudo-image of features generated from the processed pillars. A detection head [28] is then used to generate 3D bounding box for objects in point cloud data. The inference speed of PointPillars reaches 42.4 FPS, making it highly suitable for applications that demand real-time processing.
@@
PV-RCNN [29] extends SECOND by adding a keypoints branch to preserve 3D structural information. Its backbone network takes the raw point cloud data as input to extract point-wise features. PV-RCNN can achieve a higher accuracy but its inference speed is only 8.9 FPS, which falls below real-time requirements. Voxel-RCNN [26] further improves accuracy by leveraging a 3D network for feature extraction. It then applies a 2D backbone network and Region Proposal Network to BEV representation of the 3D feature volumes. Voxel-RCNN achieves an inference speed of 25.2 FPS.
CenterPoint [30] reformulates detection as a center-based approach rather than relying on anchors or dense bounding box predictions. This design improves both detection quality and inference speed, achieving 25.0 FPS on the KITTI dataset. An attention-based model, DSVT, is proposed in [31], which selectively processes only the important sparse voxel regions, improving both efficiency and scalability. With this approach, DSVT achieves both high accuracy and fast inference, delivering a runtime of 15.9 FPS on the KITTI dataset. CenterPoint and DSVT provide both voxel-based and pillar-based models. The inference speed of different models is summarized in Table I.
TABLE I: Inference Speed of Different Models on KITTI Dataset using NVIDIA RTX 2080 Ti GPU
Model FPS
SECOND 30.4
PointPillars 42.4
PV-RCNN 8.9
Voxel-RCNN 25.2
CenterPoint 25.0
DSVT 15.9
@@
While there are various 3D object detection networks that can achieve higher detection performance, we selected PointPillars for our work due to its real-time inference capability; it consistently achieves the highest FPS compared to other detection models [26]. This is a critical consideration for our safety-oriented approach, which requires the ability to process incoming point clouds in real time. While higher detection performance is desirable, it often increases computation time and resource demands, potentially hindering real-world deployment. We chose PointPillars as it can fulfill the real-time requirements of our drone detection applications.
@@
Instead of purely using the standard PointPillars network, we modify it for an air-to-air drone scenario by extending the scan space vertically. To reduce inference latency, we propose a new sparse convolution algorithm to accelerate the backbone layer, which is the most time-consuming part of PointPillars [32].
Figure 1: The PointPillars architecture consists of these three key building blocks: Pillar Feature Net, Backbone, and Detection Head.
@@ 2.2 Latency Reduction
While increasing point cloud density can improve detection accuracy, minimizing latency is crucial for ensuring real-time hazard response. The inherent sparsity of LiDAR data can be utilized to reduce latency. The study in [33] describes the spatial sparsity of feature maps and how sparsity applies to improve performance of different applications. A sparse convolutional neural network that utilizes sparse matrix multiplication is presented in [34]. An efficient sparse convolutional operation called submanifold sparse convolution is introduced in [35], in which the output will be computed if and only if the corresponding site in the input is active.
Matrix multiplication is an efficient method for implementing spatial convolution on GPU, as demonstrated by [36]. This requires transforming the input image into matrices that are suitable for fast multiplication. This transformation can be achieved by lowering the input data with duplication. The GPU implementation of the proposed method in this paper is also based on fast GPU matrix multiplication, but without data lowering.
A sparse convolution with rule book is presented in [25], where the rule book is used to map between input and output. The work presented in [37] generalizes convolution, including sparse convolution, to handle any discrete convolutions in high-dimensional spaces. A highly efficient Sparse Kernel Generator is developed in [38], which produces high performance sparse point cloud convolution kernels at a fraction of the engineering cost of the current state-of-the-art. In [39], the computational efficiency is improved by applying convolutions solely to the changed regions within successive video frames, enabling sparse feature updates to propagate throughout the network.
Most current methods implement convolution as a gathering operation, where neighboring pixels are multiplied by corresponding filter elements, and the results are summed to update the center pixel. Sparse convolution implementations typically follow this approach as well. To accelerate GPU performance, a hash table or rule book is often used to map positions between the input and output. In this paper, we propose an alternative approach to implement sparse convolution using a scattering operation. This method eliminates the need for a hash table or rule book, improving efficiency while simplifying the implementation.
@@ 2.3 Training Data Generation
The advancement of object detection research on point clouds has significantly increased the availability of datasets. Within the context of autonomous driving, there are different dataset such as KITTI [40], Waymo Open [41] and nuScenes [42] and others which are available online, and include LiDAR as part of the dataset. These datasets contain a collection of diverse urban scenes and dense position and class annotations. These datasets provide us with valuable resources for validating autonomous driving algorithms, enabling the development of more accurate and robust perception systems.
@@
LiDAR is one of the most important sensors for autonomous driving and many works have focused on the simulation of LiDAR. In [48][49], a LiDAR simulator which can augment real point clouds with virtual objects is implemented. The point cloud sweeps of real-world surroundings (usually to be used as a background) are generated from real LiDAR scanners mounted on top of a vehicle. The point cloud of a virtual object is obtained by scanning handcrafted CAD models with virtual LiDAR. These works demonstrate the importance of position and orientation of scanning objects for detection and segmentation. A LiDAR simulation approach described in [50] leverages real-world data to enhance the realism of the simulated scenes. This approach results in more diversity and realism in the simulated scenes. It also describes a learning system that models the residual physics that are not captured by graphics rendering, which further boosts the realism of the simulation. For a drone usage scenario, the study in [51] employs LiDAR simulation to evaluate the performance of detect-and-avoid systems for unmanned aerial systems (UAS).
In this work, we enhance the AI training dataset by augmenting real LiDAR data with additional synthetic point clouds generated using a Digital Twin. Unlike traditional Euclidean augmentation techniques that arbitrarily reposition objects, our approach ensures that inserted point clouds accurately reflect the LiDAR scanning pattern at their new locations. This physics-informed augmentation improves realism by preserving the sensor’s spatial and temporal characteristics, leading to more representative training data.